# Shanghai, China, 200241

### v2

@@ Abstract
In this work we uncover a connection that relates the 1-form and the 2-group symmetries of 5D SCFTs derived from geometric engineering methods to monodromies of the corresponding B-models via mirror symmetry. Viewing defects as branes wrapping relative cycles in a non-compact CY3, we find that the defect groups can be read off from the VEVs of the corresponding line operators at the leading order. Via mirror map, we find that both the 1-form and the 2-group symmetries of the SCFT are related to the monodromy at the large radius point in the B-model. Additionally, we recursively obtain closed-form expressions of instanton expansions of the VEV of Wilson lines of certain 5D theories among which some have not been obtained so far using localization methods. We further conjecture that the 2-group symmetry is given by the Mordell-Weil torsion of the universal special geometry associated to the theory, generalizing the conjecture for rank-1 theories.
@@ 1 Introduction and summary
This question can be answered when we study VEVs of the Wilson loop operators in the B-models. It was conjectured in Huang:2022hdo that the VEVs of the 5D Wilson loop operators are proportional to the complex structure parameters $u_{i}$ in the B-model of the topological string theory on the mirror manifold $W_{6}$. This correspondence can be easily verified in the toric CY3 cases where the mirror curve, a section of the B-model geometry, is precisely the Seiberg-Witten curve of $\mathcal{T}_{M_{6}}$, and the complex structure parameters $u_{i}$ are equal to the $u$-plane parameters which are the holonomies of the gauge fields. In this paper we show that, as a symmetry, the action of the one-form symmetry at the level of the partition function corresponds to the monodromy transformations for the B-model of topological strings. Furthermore, part of the monodromy actions precisely give the desired charges of the Wilson loops under one-form symmetries. We present a heuristic explanation from purely quantum field theoretical considerations of this correspondence in Section 1.1.
@@
This work is organized as follows. In Section 2.1 we start with constructing Wilson line operators from M2-brane wrapping modes in M-theory compactification then write their VEVs as functions of the Kähler moduli parameters of the internal geometry. In Section 2.2 we will discuss the relation between the 1-form symmetry of the 5D theory which is $\mathcal{T}_{M_{6}}$ on $S^{1}$ and the shift of the $B_{2}$ field in the corresponding IIA picture. In general, for a non-compact CY3, the dimension $b_{2}$ of the curve classes is larger than the dimension $b_{4}$ of the compact surface classes, providing additional Kähler parameters in the A-model and additional complex structure parameters in the B-model. These additional parameters represent flavors masses and instanton counting parameters on the extended Coulomb branch of $\mathcal{T}_{M_{6}}$. So in general, the actions of the monodromy group are not always the one-form symmetric actions, but they are the actions of the 2-group symmetry as discussed in Section 2.3. We will present our main result in Section 3.1 where we relate the monodromy of the B-model at the large radius point with the 1-form and the 2-group symmetries of the 5D theory. In Section 3.2 and 3.3 we provide concrete examples to test our statements. As a byproduct, we are able to obtain (order-by-order) closed-form results of the instanton expansions of certain theories via the application of mirror symmetry methods, which have not been calculated before via the traditional localization techniques. In Section 4.1 we will discuss some of the physical consequence of our calculation in Section 2 and 3 and propose in Section 4.2 a conjecture relating the topological data of the special geometry of the theory and its 2-group symmetry, generalizing the relevant discussions in Closset:Uplane ; Cecotti:2024jbt . In Section 5 we summarize our results and discuss potential applications and generalizations of the method developed in this work.
1.1 Motivation from quantum field theory considerations
@@
In this section, we discuss some general aspects of the action of the symmetry operator upon a defect using the path integral formalism of quantum field theory. The discussion in this section is intended to motivate the rest of the work. We will make precise the heuristic arguments in this section in the main part of this work.
@@
We will focus on abelian higher form symmetries of certain quantum field theory $\mathcal{T}$. It is known that the action of a topological symmetry operator $\mathcal{O}$ supported on $\Sigma$ upon a charged defect $W$ supported on $\ell$ is Gaiotto:GenSymm :
@@
In the above equation, the dependence of $W(\ell)$ on $\phi$ is omitted for simplicity. The absorption in (7) can be done, e.g. for an abelian global symmetry with conserved Noether current $j$ in which case $\mathcal{O}(\Sigma)=e^{i\int_{\Sigma}\star j}$ with closed $\Sigma$. Therefore we have $\mathcal{O}(\Sigma)=e^{i\int_{M^{d}}A_{\Sigma}\wedge\star j}$ where $A_{\Sigma}$ is the Poincaré dual of $\Sigma$ in $M^{d}$. Thus we arrive at the new action $S^{\prime}=S+\int_{M^{d}}A_{\Sigma}\wedge\star j$. Suppose that there exists a change of variables $\phi\rightarrow\phi^{\prime}$ such that the functional form of $S$ is recovered while the dependence of $W(\ell)$ on $\phi$ remains unchanged. Therefore we have:
@@
The existence of such a field redefinition can be seen by considering the symmetry action on a trivial defect $W=1$ as follows:
$\mathcal{Z}=\int\mathcal{D}\phi\ e^{iS^{\prime}[\phi]}=\int\mathcal{D}\phi\ e^{iS[\phi]}\stackrel{{\scriptstyle\text{rename dummy variable}}}{{=}}\int\mathcal{D}\phi^{\prime}\ e^{iS[\phi^{\prime}]}$ (9)
where the first equality comes from the RHS of (7) and the second equality is nothing but the definition of $\mathcal{Z}$.
@@
We will see in the following sections that the above heuristic arguments from a purely quantum-field-theoretical point of view can be made very precise. Indeed, we will show that $z$ is the complex structure moduli of the mirror of the Calabi-Yau threefold on which we compactify M-theory to geometrically engineer theory $\mathcal{T}$. In particular, the monodromy exhibited by the toy model (13) will be made precise in Section 2 and 3 using mirror symmetry and indeed we will find that $\langle W(\ell)\rangle=z^{-1/n}$ at the leading order of the mirror map.
@@ 2 Defect lines from geometric engineering and the mirror
In this Section we discuss some basic properties of Wilson lines in 5D theory from various angles. In particular, we are interested in the Wilson lines that are infinitely massive defects, which are extended charged objects under the 1-form symmetry of the 5D theory Gaiotto:2014ina . We study the geometrically engineered 5D theory via M-theory on a non-compact CY3 $M_{6}$, which we denote by $\mathcal{T}_{M_{6}}$. We will further reduce $\mathcal{T}_{M_{6}}$ on $S^{1}$ and discuss the relevant properties of the resulting 4D KK theory.
@@ 2.1 Defect lines from wrapped M2-brane branes
where in here and in the following the (relative) homology groups are taken with coefficient group to be $\mathbb{Z}$ unless otherwise specified. In this work we will always assume $M^{5}=\mathbb{R}^{4}\times S^{1}$ and set $\ell=S^{1}$. In this sense we are actually studying the $S^{1}$-compactified 5D theories, or equivalently, the 4D $\mathcal{N}=2$ Kaluza-Klein type (KK) theories. The classical action of such an M2-brane is ×$\int_{\ell\times\Sigma}C_{3}$ plus a volume form, the integration of $C_{3}$ provides the antisymmetric $B_{2}$ field or in short $B$-field in the resulting IIA theory while the integration over the volume form gives the volume of $\Sigma$, so the total contribution gives provides a complexified Kähler parameter $t_{\Sigma}$. The VEV of the Wilson loop operator counts the numbers of such M2-branes on $\Sigma^{\prime}\in H_{2}(M_{6},\partial M_{6};\mathbb{Z})$ which has degree one component of $\Sigma$, therefore we have:
@@
where $t_{i}$ are the complexified Kähler parameters of $M_{6}$. The leading coefficient in (16) is always one because in the infinity massive limit $R\rightarrow\infty$, only single non-dynamic M2-brane contributes.
@@
In the infinitely massive limit $t_{\Sigma}\rightarrow\infty$, the massive quark is decoupled from the system, thus the VEV of the Wilson loop (16) is zero. To make sense of the calculation, the concept compact representative was introduced in Bhardwaj:Higher_form_5D6D ; Hubner:GenSymm_EFib . Recall that the 11D three-form gauge field $C_{3}$ has a decomposition $C_{3}=A\wedge\omega^{\Sigma}$ on $\Sigma$, for $\omega^{\Sigma}\in H_{\text{cpt}}^{2}(M_{6},\partial M_{6})$ such that the pairing $\kappa_{R}:=\int_{\Sigma}\omega^{\Sigma}$ is non-trivial, which is essentially the volume of $\Sigma$ proportional to the cut-off $R$  [^2] Actually, as a cone over a 1-cycle $\gamma\subset\partial M_{6}$, $\Sigma$ depends on $R^{d}$ for certain $d$. But this power is not important for our purpose.. If we take the limit $R\rightarrow\infty$, the volume of $\Sigma$, which is proportional to the mass of the massive quark, goes to infinity. Subtracting this infinity contribution we find a finite result for the VEV of the Wilson loop operator, which can be computed by introducing the compact representative
@@
where $z$ is the mirror of $t$ which is the Kähler parameter of the compact representative corresponding to the generator of $Z(SU(N))\cong\mathbb{Z}_{N}$. It is immediate to see that the form of $\langle W\rangle$ matches our toy model (13), therefore by our heuristic argument in Section 1.1 the theory has $\mathbb{Z}_{N}$ 1-form symmetry which is indeed true for pure $SU(N)$ theory. We conclude that the application of mirror map upon the VEV of Wilson loop backs up our heuristic derivation in 1.1. We will sharpen this observation further in Section 3 with more detailed discussions and concrete examples and see it encodes not only the 1-form symmetry but also the 2-group symmetry of $\mathcal{T}_{M_{6}}$.
@@ 2.3 A subtlety that signals 2-group symmetry
For instance, let us consider a generic 5D KK theory arises from circle compactification of a 6D SCFT, where the 6D theory is obtained from the F-theory compactified on the elliptically fibered Calabi-Yau threefold $M_{6}$ over a non-compact base. Followed by the discussion in Cvetic:0form_1form_2group , let $\Delta=\cup_{i}\Delta_{i}\subset M_{6}$ where $\Delta_{i}$ is the $F_{i}$-flavor brane locus and we study the geometry of $\partial M_{6}^{\circ}\cap T(\Delta_{i})$ illustrated in Figure 3 where $\partial M_{6}^{\circ}:=\partial M_{6}\backslash\Delta$ and $T(\Delta_{i})$ is a tubular neighborhood of $\Delta_{i}$.
@@
In this section, we provide a detailed discussion on how to calculate the VEVs for 5D Wilson loops in the topological string B-model, as well as the method for determining 1-form and 2-group symmetries in (26). We refer the reader to Clader:2018yyu for a review on the topological string B-models. In section 3.1 we present our main statement that the 1-form and the 2-group symmetries of the 5D theory $\mathcal{T}_{M_{6}}$ can be identified with certain reparametrization groups of the mirror geometry. In Section 3.2 and 3.3 we present concrete examples to illustrate and test our main claim.
@@
then the vanishing loci of $\mathcal{H}(e^{2\pi i\,\lambda}u,e^{2\pi i\,\kappa}m;Y)$ and $\mathcal{H}(u,m;Y)$ are the same hypersurface and they describe the same physics system. See Lerche:1989cs ; Lerche:1991wm for related discussions.
@@
It has been proposed in Huang:2022hdo , the VEVs of the half-BPS Wilson loop operators in the 5D supersymmetric quantum field theory on $S^{1}$ can be calculated in its corresponding topological string B-model; for the 5D $SU(N)$ theory, at genus zero, we have  [^4] As we will see in Section 3.2 and 3.3, a more generic form of the Wilson loop is $\left\langle W_{\mathbf{R}_{i}}\right\rangle=u_{i}(1+f(z))$, where $f(z)$ is usually a finite polynomial of $z_{i}$. For all the pure gauge theories we will consider, we find $\left\langle W_{\mathbf{R}_{i}}\right\rangle=u_{i}+\sum_{j}P_{j}u_{j}$, where the summation is over all $u_{j}$’s that have the same charge as $u_{i}$ under the one-form symmetry. $P_{j}$ is a constant that only depends on mass parameters.
@@
It is natural to expect that when there are flavor branes, in particular their existence breaks the 1-form symmetry, (40) defines VEV for a “naive” line defect, which carries charges under $\mathcal{G}_{um}\cong\mathcal{E}$. As mentioned in Section 2.1, to calculated the VEV of a Wilson loop operator from M2 branes on the relate 2-cycle $\Sigma$, we need to specify the compact representative of $\Sigma$. However, ambiguities arise from the non-uniqueness of this compact representative. Different choices of compact representatives correspond to different choices of non-compact surfaces and can be related through reparametrization of the moduli parameters $u_{i}$. This results in the fact that (40) may differ by a factor that depends only on the mass parameters. As we will demonstrate in Section 3.2 and 3.3, the charge of (40) under $\mathcal{E}$ remains invariant under reparametrization of $u_{i}$, ensuring that (40) is a consistent observable.

### v3

@@ Abstract
In this work we discuss a connection that relates the 1-form and the 2-group symmetries of 5D SCFTs derived from geometric engineering methods to monodromies of the corresponding B-models via mirror symmetry. Viewing defects as branes wrapping relative cycles in a non-compact CY3, we find that the defect groups can be read off from the VEVs of the corresponding line operators at the leading order. Via mirror map, we find that both the 1-form and the 2-group symmetries of the SCFT are related to the monodromy at the large radius point in the B-model. Additionally, we recursively obtain closed-form expressions of instanton expansions of the VEV of Wilson lines of certain 5D theories among which some have not been obtained so far using localization methods. We further conjecture that the 2-group symmetry is given by the Mordell-Weil torsion of the universal special geometry associated to the theory, generalizing the conjecture for rank-1 theories.
@@ 1 Introduction and summary
This question can be answered when we study VEVs of the Wilson loop operators in the B-models. It was conjectured in Huang:2022hdo that the VEVs of the 5D Wilson loop operators are proportional to the complex structure parameters $u_{i}$ in the B-model of the topological string theory on the mirror manifold $W_{6}$. This correspondence can be easily verified in the toric CY3 cases where the mirror curve, a section of the B-model geometry, is precisely the Seiberg-Witten curve of $\mathcal{T}_{M_{6}}$, and the complex structure parameters $u_{i}$ are equal to the $u$-plane parameters which are the holonomies of the gauge fields. In this paper we show that, as a symmetry, the action of the one-form symmetry at the level of the partition function corresponds to the monodromy transformations for the B-model of topological strings. Furthermore, part of the monodromy actions precisely gives the desired charges of the Wilson loops under one-form symmetries. We present a heuristic explanation from purely quantum field theoretical considerations of this correspondence in Section 2.
@@
This work is organized as follows. In Section 2 we relate the generalized symmetries of a QFT to certain monodromies in its moduli space via a somewhat heuristic discussion from pure QFT perspective, which will be made concrete and precise in subsequent sections via mirror symmetry. In Section 3.1 we construct Wilson line operators from M2-brane wrapping modes in M-theory compactification then write their VEVs as functions of the Kähler moduli parameters of the internal geometry. In Section 3.2 we will discuss the relation between the 1-form symmetry of the 5D theory which is $\mathcal{T}_{M_{6}}$ on $S^{1}$ and the shift of the $B_{2}$ field in the corresponding IIA picture. In general, for a non-compact CY3, the dimension $b_{2}$ of the curve classes is larger than the dimension $b_{4}$ of the compact surface classes, providing additional Kähler parameters in the A-model and additional complex structure parameters in the B-model. These additional parameters represent flavors masses and instanton counting parameters on the extended Coulomb branch of $\mathcal{T}_{M_{6}}$. So in general, the actions of the monodromy group are not always the one-form symmetric actions, but they are the actions of the 2-group symmetry as discussed in Section 3.3. We will present our main result in Section 4.1 where we relate the monodromy of the B-model at the large radius point with the 1-form and the 2-group symmetries of the 5D theory. In Section 4.2 and 4.3 we provide concrete examples to test our statements. As a byproduct, we are able to obtain (order-by-order) closed-form results of the instanton expansions of certain theories via the application of mirror symmetry methods, which have not been calculated before via the traditional localization techniques. In Section 5.1 we will discuss some of the physical consequence of our calculation in Section 3 and 4 and propose in Section 5.2 a conjecture relating the topological data of the special geometry of the theory and its 2-group symmetry, generalizing the relevant discussions in Closset:Uplane ; Cecotti:2024jbt . In Section 6 we summarize our results and discuss potential applications and generalizations of the method developed in this work.
2 Motivation from quantum field theory considerations
@@
In this section, we review some general aspects of the action of the symmetry operator upon a defect using the path integral formalism of quantum field theory (see e.g. Gomes:review ; Bhardwaj:lecture ; Schafer-Nameki:ICTP ; Jia:2025jmn ). The discussion in this section is intended to motivate the rest of the work. In particular, we aim to illustrate the relationship between certain monodromy actions in the moduli space of a QFT and its generalized symmetries  [^1] More precisely, the generalized symmetries to be discussed are actually the symmetries at the SCFT point in the moduli space of the QFT, and we will see in Section 3 that the monodromy is the one around the point of large complex structure limit in the moduli space of the mirror geometry.. We will make the following somewhat heuristic arguments precise in the subsequent sections.
@@
We will focus on abelian higher form symmetries of certain quantum field theory $\mathcal{T}$ in spacetime $M_{d}$. It is known that the action of a topological symmetry operator $\mathcal{O}$ supported on $\Sigma$ upon a charged defect $W$ supported on $\ell$ is Gaiotto:GenSymm :
@@
In the above equation, the dependence of $W(\ell)$ on $\phi$ is omitted for simplicity. The absorption in (7) can be done, e.g. for an abelian global symmetry with conserved Noether current $j$ in which case $\mathcal{O}(\Sigma)=e^{i\int_{\Sigma}\star j}=e^{i\int_{M^{d}}A_{\Sigma}\wedge\star j}$ with closed $\Sigma$ and $A_{\Sigma}$ is the Poincaré dual of $\Sigma$ in $M^{d}$. Thus we arrive at the new action $S^{\prime}=S+\int_{M^{d}}A_{\Sigma}\wedge\star j$. Suppose that there exists a change of variables $\phi\rightarrow\phi^{\prime}$ such that the functional form of $S$ is recovered while the dependence of $W(\ell)$ on $\phi$ remains unchanged, we have:
@@
To see the existence of such a change of variables, let us look at $U(1)$ gauge theory with electric 1-form symmetry as a concrete example. In this case the topological operator is Bhardwaj:lecture :
$\mathcal{O}(\Sigma)=\exp\left(i\alpha\int_{\Sigma}\star F\right),\ g=e^{i\alpha}$ (9)
for a $(d-2)$-dimensional closed surface $\Sigma$, which can be further rewritten as:
$\mathcal{O}(\Sigma)=\exp\left(i\alpha\int_{M_{d}}\delta_{\Sigma}\wedge\star F\right)$ (10)
for a 2-form $\delta_{\Sigma}$ with delta function support on $\Sigma$. To obtain the absorption (7), we consider the shift $A\rightarrow A-\frac{1}{2}\frac{\alpha}{2\pi}d\varphi$, where $\varphi$ parametrizes a small circle linking $\Sigma$ in $M_{d}$ and $d(d\varphi)=2\pi\delta_{\Sigma}$ Beasley:2009mb . For such shift, we have:
$S[A]\rightarrow S^{\prime}[A]=S[A]-\alpha\int_{M_{d}}\delta_{\Sigma}\wedge\star F+\frac{\alpha^{2}}{4}\int_{M_{d}}\delta_{\Sigma}\wedge\star\delta_{\Sigma}$ (11)
where the second term is exactly what is needed to absorb the insertion of $\mathcal{O}(\Sigma)$, and the last term represents a formally divergent “self energy” of $\Sigma$ to be regularized by adding a counter term  [^2] In 3D for Chern-Simons theory this would be the self-linking of a knot, which can be regularized by the point-splitting regularization Witten:1988hf . Here in principle we care only about the correlator $\langle\mathcal{O}_{\Sigma}W(\ell)\rangle$ and the self-linking, or self energy, of the operators are unimportant for our purpose.. The shift $A\rightarrow A-\frac{1}{2}\frac{\alpha}{2\pi}d\varphi$, however, does not change the functional dependence of $W(\ell)=e^{i\oint_{\ell}A}$ on $A$ at all since by definition $\ell\cap\Sigma=\emptyset$ hence any term proprotional to $\delta_{\Sigma}$ integrates to zero over $\ell$. Therefore, in this concrete example, apart from a divergent term to be regularized, the shift $A\rightarrow A-\frac{1}{2}\frac{\alpha}{2\pi}d\varphi$ is the desired field redefinition for (8) to hold.
@@
The lesson we learn from the discussions in this section is that the higher form symmetry of a QFT, which traditionally is viewed as a property of the QFT itself by manipulating its local fields as in (8) and (13), can actually be related to monodromies in its moduli space. We will see in the following sections that the above heuristic arguments from a purely quantum-field-theoretical point of view can be made very precise. Indeed, we will show that $z$ is the complex structure moduli of the mirror of the Calabi-Yau threefold on which we compactify M-theory to geometrically engineer theory $\mathcal{T}$. In particular, the monodromy exhibited by the toy model (15) will be made precise in Section 3 and 4 using mirror symmetry and indeed we will find that $\langle W(\ell)\rangle=z^{-1/n}$ at the leading order of the mirror map.
@@ 3 Defect lines from geometric engineering and the mirror
In this Section we discuss some basic properties of Wilson lines in 5D theory from various angles. In particular, we focus on the geometrically engineered 5D theory via M-theory on a non-compact CY3 $M_{6}$, which we denote by $\mathcal{T}_{M_{6}}$, and study the Wilson lines that are obtained from M2-brane wrapping non-compact cycles that are extended charged objects under the 1-form symmetry of the 5D theory Gaiotto:2014ina . We will further reduce $\mathcal{T}_{M_{6}}$ on $S^{1}$ and discuss the relevant properties of the resulting 4D KK theory.
@@ 3.1 Defect lines from wrapped M2-brane branes
where in here and in the following the (relative) homology groups are taken with coefficient group to be $\mathbb{Z}$ unless otherwise specified. In this work we will always assume $M^{5}=\mathbb{R}^{4}\times S^{1}$ and set $\ell=S^{1}$. In this sense we are actually studying the $S^{1}$-compactified 5D theories, or equivalently, the 4D $\mathcal{N}=2$ Kaluza-Klein type (KK) theories. The classical action of such an M2-brane is ×$\int_{\ell\times\Sigma}C_{3}$ plus a volume form, the integration of $C_{3}$ provides the antisymmetric $B_{2}$ field or in short $B$-field in the resulting IIA theory while the integration over the volume form gives the volume of $\Sigma$, so the total contribution gives provides a complexified Kähler parameter $t_{\Sigma}$. The VEV of the Wilson loop operator gives the number of such M2-branes on $\Sigma^{\prime}\in H_{2}(M_{6},\partial M_{6};\mathbb{Z})$ which has degree one component of $\Sigma$, therefore we have:
@@
where $t_{i}$ are the complexified Kähler parameters of $M_{6}$. The leading coefficient in (18) is always one because in the limit $R\rightarrow\infty$, only single non-dynamic M2-brane contributes.
@@
In the limit $t_{\Sigma}\rightarrow\infty$, the massive quark is decoupled from the system, thus the VEV of the Wilson loop (18) is zero. To make sense of the calculation, the concept compact representative was introduced in Bhardwaj:Higher_form_5D6D ; Hubner:GenSymm_EFib . Recall that the 11D three-form gauge field $C_{3}$ has a decomposition $C_{3}=A\wedge\omega^{\Sigma}$ on $\Sigma$, for $\omega^{\Sigma}\in H_{\text{cpt}}^{2}(M_{6},\partial M_{6})$ such that the pairing $\kappa_{R}:=\int_{\Sigma}\omega^{\Sigma}$ is non-trivial, which is essentially the volume of $\Sigma$ proportional to the cut-off $R$  [^4] Actually, as a cone over a 1-cycle $\gamma\subset\partial M_{6}$, $\Sigma$ depends on $R^{d}$ for certain $d$. But this power is not important for our purpose.. If we take the limit $R\rightarrow\infty$, the volume of $\Sigma$, which is proportional to the mass of the massive quark, goes to infinity. Subtracting this infinity contribution we find a finite result for the VEV of the Wilson loop operator, which can be computed by introducing the compact representative
@@
where $z$ is the mirror of $t$ which is the Kähler parameter of the compact representative corresponding to the generator of $Z(SU(N))\cong\mathbb{Z}_{N}$. It is immediate to see that the form of $\langle W\rangle$ matches our toy model (15), therefore by our heuristic argument in Section 2 the theory has $\mathbb{Z}_{N}$ 1-form symmetry which is indeed true for pure $SU(N)$ theory. We conclude that the application of mirror map upon the VEV of Wilson loop backs up our heuristic derivation in 2. We will sharpen this observation further in Section 4 with more detailed discussions and concrete examples and see it encodes not only the 1-form symmetry but also the 2-group symmetry of $\mathcal{T}_{M_{6}}$.
@@ 3.3 A subtlety that signals 2-group symmetry
For instance, let us consider a generic 5D KK theory arises from circle compactification of a 6D SCFT, where the 6D theory is obtained from the F-theory compactified on the elliptically fibered Calabi-Yau threefold $M_{6}$ over a non-compact base. Following the discussion in Cvetic:0form_1form_2group , let $\Delta=\cup_{i}\Delta_{i}\subset M_{6}$ where $\Delta_{i}$ is the $F_{i}$-flavor brane locus and we study the geometry of $\partial M_{6}^{\circ}\cap T(\Delta_{i})$ illustrated in Figure 3 where $\partial M_{6}^{\circ}:=\partial M_{6}\backslash\Delta$ and $T(\Delta_{i})$ is a tubular neighborhood of $\Delta_{i}$.
@@
In this section, we provide a detailed discussion on how to calculate the VEVs for 5D Wilson loops in the topological string B-model, as well as the method for determining 1-form and 2-group symmetries in (28). We refer the reader to Clader:2018yyu for a review on the topological string B-models. In section 4.1 we present our main statement that the 1-form and the 2-group symmetries of the 5D theory $\mathcal{T}_{M_{6}}$ can be identified with certain reparametrization groups of the mirror geometry. In Section 4.2 and 4.3 we present concrete examples to illustrate and test our main claim.
@@
then the vanishing loci of $\mathcal{H}(e^{2\pi i\,\lambda}u,e^{2\pi i\,\kappa}m;Y)$ and $\mathcal{H}(u,m;Y)$ are the same hypersurface and they describe the same physical system. See Lerche:1989cs ; Lerche:1991wm for related discussions.
@@
It has been proposed in Huang:2022hdo , the VEVs of the half-BPS Wilson loop operators in the 5D supersymmetric quantum field theory on $S^{1}$ can be calculated in its corresponding topological string B-model; for the 5D $SU(N)$ theory, at genus zero, we have  [^6] As we will see in Section 4.2 and 4.3, a more generic form of the Wilson loop is $\left\langle W_{\mathbf{R}_{i}}\right\rangle=u_{i}(1+f(z))$, where $f(z)$ is usually a finite polynomial of $z_{i}$. For all the pure gauge theories we will consider, we find $\left\langle W_{\mathbf{R}_{i}}\right\rangle=u_{i}+\sum_{j}P_{j}u_{j}$, where the summation is over all $u_{j}$’s that have the same charge as $u_{i}$ under the one-form symmetry. $P_{j}$ is a constant that only depends on mass parameters.
@@
It is natural to expect that when there are flavor branes, in particular their existence breaks the 1-form symmetry, (42) defines VEV for a “naive” line defect, which carries charges under $\mathcal{G}_{um}\cong\mathcal{E}$. As mentioned in Section 3.1, to calculated the VEV of a Wilson loop operator from M2 branes on the relate 2-cycle $\Sigma$, we need to specify the compact representative of $\Sigma$. However, ambiguities arise from the non-uniqueness of this compact representative. Different choices of compact representatives correspond to different choices of non-compact surfaces and can be related through reparametrization of the moduli parameters $u_{i}$. This results in the fact that (42) may differ by a factor that depends only on the mass parameters. As we will demonstrate in Section 4.2 and 4.3, the charge of (42) under $\mathcal{E}$ remains invariant under reparametrization of $u_{i}$, ensuring that (42) is a consistent observable.